Please Help me with this word problem.

Six grapefruit cost as much as a dozen oranges. The cost of a dozen grapefruit and two dozen oranges is $12.00. How much does one grapefruit cost?

Thanks.

Answer by checkley71(8403)

6GRAPEFRUIT=12ORANGES
12G+24O=12.00 THEREFORE G=12/6 ORANGES OR G=2 ORANGES THEREFORE
12*2 0RANGES+24 ORANGES=12.00 OR 24+24=12.00 OR 48=12.00 OR EACH ORANGE =$.25
THEN 12G+24*.25=12.00 OR 12G+6.00=12 OR 12G=6.00 OR G=6.00/12 OR GRAPEFRUIT=.50
PROOF 6 GRAPEFRUIT @ .50 $3.00=12 ORANGES @ .25 OR $3.00 AND
12*.50+24*.25=$12.00 OR $6.00+$6.00=$12.00